Grade 9Algebra

if a=xy^p-1 , b=xy^q-1 and c=xy^r-1, prove that a^q-r b^r-p c^p-q = 1

harshita
8 Years agoGrade 9

2 Answers

Deepak Kumar Shringi
8 Years ago
is it power of xy or power of y only..kindly upload a image that will be more helpful
Arun
8 Years ago
hence provedLHS = x^{0}y^{pq-pr-q+r+qr-pq-r+p+pr-qr-p+q} = x^0y^0 = 1 = RHSLHS = x^{q-r+r-p+p-q}y^{(p-1)(q-r) + (q-1)(r-p) + (r-1)(p-q)}so to do that just substitute the values of a b and cwe geta^{q-r}b^{r-p}c^{p-q} = 1you would have to prove
 
 
Regards
Arun (askIITians forum expert)